# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

### v1

@@ Abstract
Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods. While recent advances allowed LLMs to interact with online tools, solve robotics tasks and many more, long range reasoning tasks remain a problem for LLMs. Existing methods to address this issue are very resource intensive and require additional data or human crafted rules, instead, we propose a simple method based on few-shot in-context learning alone to enhance ‘chain-of-thought’ with state-tracking for planning and acting with LLMs. We show that our method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution. We also demonstrate that our enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in number of steps required to solve a task. We show that our method works across a variety of LLMs for both API-based and open source ones. Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ helps state-tracking accuracy, while a json-structure harms overall performance. We open-source our code and annotations at https://github.com/ai-nikolai/StateAct.
@@ 1 Introduction
Using the in-build world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020) for example. of large language models (LLMs), such as GPT-3, Mixtral, Gemini Brown et al. (2020); Jiang et al. (2024); Anil et al. (2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al. (2024). These range from solving (simulated) robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks such as online shopping Yao et al. (2023a) and navigating operating systems Liu et al. (2023), to playing a variety of games Côté et al. (2019); Liu et al. (2023).
While LLMs are increasingly successful in solving AI agent tasks, Li et al. (2023b) show that LLMs struggle with long range understanding and Coelho et al. (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem. Existing efforts to address these issues are resource intensive, Wu et al. (2024) required human expert annotations of rules, Sun et al. (2023) a code execution environment with carefully crafted code-based prompts, Fu et al. (2024) use additional training data together with retrieval augmented generation (RAG) to help the AI agent.
To overcome these challenges, we introduce “StateAct” a novel LLM agent based on few-shot in-context learning that tracks the goal by ‘reminding’ the agent of the goal and explicitly keeps track of the agent’s state (such as location and inventory). We also utilise ‘chain-of-thought’ (i.e. explicit reasoning Wei et al. (2023)) as an optional addition. At the core of the method lies a simple annotation with goal-, state- and reasoning- traces of the few-shot examples, that we call ‘chain-of-states’.
Our method establishes a new state-of-the-art for Alfworld Shridhar et al. (2021) for few-shot in-context learning based methods, outperforming the previous best method by 14% and even outperforms methods that use external tools by 2.5%. Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource intensive and requires additional data and expert human annotations. Comparing to in-context learning methods, we also demonstrate that our method is both more efficient in terms of number of steps to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
@@ 2 Background
AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018). With the dawn of LLMs, works such as Li et al. (2022); Nottingham et al. (2023) combined LLMs and RL and trained additional policies or value functions to make predictions.
@@ 2.1 In-context learning approaches
Huang et al. (2022a, b); Singh et al. (2022) were among the first to use LLMs directly to act in an interactive environment, their method produces agent actions as output after receiving environment observations as input.
ReAct Yao et al. (2023b) took this work further by combining ‘acting‘ Huang et al. (2022a) and ‘chain-of-thought‘ Wei et al. (2023). ReAct establishes state-of-the-art for in-context learning only based approaches and while it is a very scalable method, the performance (i.e. success rate) is still limited.
ExpeL Zhao et al. (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training. At inference time they look-up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference. They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
@@ 2.1.1 In-context learning and additional tools
The current state-of-the-art for in-context learning based approaches in combination with additional tools is AdaPlanner Sun et al. (2023). They introduce a code-based prompt Li et al. (2023a) and use code-execution as an additional tool to execute the LLM generations to feed them back into the next prompt. The shortcoming of AdaPlanner is that it requires very complex human crafted prompts that are hard to scale to new environments as well as the additional step of requiring code-execution.
@@ 2.2 State tracking in LLM-based agents
Concurrent work to ours, AutoGuide Fu et al. (2024), uses ReAct as the base agent and additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent, they then use a type of retrieval augmented generation (RAG) process to guide the decision making process. They embed the current observation as a summary (this is what they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action. Using training data and LLM-based RAG they achieve 79.1% on top of a ReAct agent on Alfworld. Their training and RAG approach could be used complementary to our StateAct LLM agent [^2] The cost of using RAG at every generation step could be significantly higher than our method, however..
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task, without training data. Their method differs to ours two-fold. Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and finally a prediction of a program). Secondly, their system involves execution of actual programs. Our method on the hand requires a straight-forward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action and we do not require program execution.
@@ 2.2.1 Fine-tuning approaches
2.4 Joined rule and LLM based agents
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human-expert crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘experts’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al. (2021)).
@@ 2.5 Approaches with multiple trials
Works such as Reflexion Shinn et al. (2023) or work by Sun et al. (2023); Yang et al. (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score. While the score necessarily goes up (as only the failed environments are re-run) we do not dive into this, as this does not compare to the actual agent performance and ‘reflection‘ can be used on top of most LLM-based agents (e.g. Reflexion uses ReAct as the base agent).
@@ 3 Method
StateAct is an LLM-based AI agent that works on top of pre-trained large language models. StateAct takes the textual ‘observation’ from the environment and after a single call to the pre-trained LLM returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current ‘interaction trace’ up to the current step in the environment, see Figure 1(b). An ‘interaction trace’ consists of alternating ‘observations’ from the environment and desired (or actual) ‘outputs’ from the LLM. In the case of StateAct the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The ‘action’ is then extracted and passed to the environment to produce the next observation, see Figure 2. For example a sample of the trace for step n in Alfworld:
@@
In the above example, the agent produces the action ‘take tomato 1’, in response to the observation ‘On the countertop 1, you see a tomato 1’.
@@ 3.1 Goal-reminders and state- tracking using LLMs
In order to make StateAct more precise, we introduce simple notation. By denoting $\pi$ as the policy of an AI agent, in the standard case at time step $t$ the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
@@
Where $a_{t}$ is the action produced by the agent at step t and $o_{t}$ is the observation produced by the environment at step t after receiving action $a_{t}$ as input. Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
For our case we need to enhance the policy to incorporate the ‘state’. Similar to previous work Yao et al. (2023b) we introduce the context vector, $c_{t}$. The context vector contains the action, as well as the other additional predictions of the agent, i.e. $c_{t}=(g_{0},s_{t},r_{t},a_{t})$. Where $g_{0}$ is the goal and always remains the same (for a given environment) and uses the goal extracted from $o_{0}$, $s_{t}$ represents the predicted state at time step $t$, $r_{t}$ represents ‘chain-of-thought’ style ‘reasoning’ at time step $t$, and $a_{t}$ represents the action at time step $t$, as before. The new policy $\pi$ then becomes:
@@
In our case the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
@@ 4 Experimental setup
Our aim is to study long-range acting, planning and reasoning capabilities of LLM-based agents. To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models. Specifically, we use Alfworld Shridhar et al. (2021), a household robotics environment, and Webshop Yao et al. (2023a), an online shopping environment, as simulated environments. In line with previous work we use OpenAI’s [^3] https://openai.com, last accessed June 2024. gpt-3.5-turbo-1106. Additionally, we evaluate on additional LLMs and validate our results.
@@ 4.1 Alfworld
Alfworld Shridhar et al. (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language based AI models, see Figure 3. Alfworld has a total of 135 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step the environment provides a textual observation or feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019). See Appendix A for a complete list of commands and details on environments.
@@ 4.1.1 Alfworld correction
In our research we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way and using only “in” or only “on” produces a failed command. We observed this issue with LLMs on this environment and we propose a simple fix for it. We map: 1. “put <object> in <place>” and 2. “put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
Methods such as AdaPlanner Sun et al. (2023) have avoided this issue because they use code-based prompts and regex parsers. However, methods such as ReAct Yao et al. (2023b) and ExpeL Zhao et al. (2023) have been affected, lowering their potential performance. In our work, we also report the results for ReAct using corrections.
@@ 4.2 Webshop
Figure 4: An example textual interaction in Webshop (right) and corresponding e-commerce website rendering (left).
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results and select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product. See Appendix B for more details. In total the test set consists of 500 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of 15 steps and two commands: 1. search[<query>], 2. click[<button>].
@@ 4.3 In context learning
Since ReAct Yao et al. (2023b) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024), we use the same few-shot ‘interaction traces’ as ReAct. The main reason is to have a fair comparison and isolate additional effect - such as performance change from different in-context examples.
In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average each ReAct example ranges from 352 words to 591 words (590 token to 935 tokens). For our study we reuse the observations, thoughts and actions and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation we discovered minor errors in the ReAct prompts and fixed them as well. We release all our annotations with our code release. In comparison, AdaPlanner Sun et al. (2023), uses a different code based approach and the prompt has 1104 words (2015 tokens) on average.
We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop for all our experiments.

### v2

@@ Abstract
Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods. While recent advances allowed LLMs to interact with online tools and solve robotics tasks, long range reasoning tasks remain a problem for LLMs. Existing methods to address this issue are very resource-intensive and require additional data or human crafted rules; instead, we propose StateAct, a simple method based on few-shot in-context learning alone. StateAct enhances ‘chain-of-thought’ with state- and goal-tracking to help with planning and acting with LLMs. The method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution. We also demonstrate that StateAct with enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in the number of steps required to solve a task. We show that StateAct works across a variety of API-based and open source LLMs. Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ help state-tracking accuracy, while a json-structure harms overall performance. We open-source our code and annotations at https://github.com/ai-nikolai/stateact.
@@ 1 Introduction
Using the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020), for example. of large language models (LLMs), such as GPT-3, Mixtral and Gemini Brown et al. (2020); Jiang et al. (2024); Anil et al. (2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al. (2024). These range from solving simulated robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks, such as online shopping Yao et al. (2023a) and navigating operating systems Liu et al. (2023), to playing a variety of games Côté et al. (2019); Liu et al. (2023).
While LLMs are increasingly successful in solving AI agent tasks, Li et al. (2023b) show that LLMs struggle with long range understanding and Coelho et al. (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem. Existing efforts to address these issues are resource-intensive: Wu et al. (2024) required human expert annotations of rules, Sun et al. (2023) required a code execution environment with carefully crafted code-based prompts, and Fu et al. (2024) used additional training data together with retrieval augmented generation (RAG) to help the AI agent.
To overcome these challenges, we introduce “StateAct”, a novel LLM-agent based on few-shot in-context learning. The method tracks the goal by ‘reminding’ the agent of the goal. StateAct, also, explicitly keeps track of the agent’s state (such as location and inventory). Additionally, we utilise ‘chain-of-thought’ (i.e. explicit reasoning Wei et al. (2023)) as an optional addition. At the core of the method lies a straightforward annotation of the few-shot examples, specifically with goal-, state- and reasoning- traces, which we call ‘chain-of-states’.
Our method establishes a new state-of-the-art for Alfworld Shridhar et al. (2021) for few-shot in-context learning-based methods, outperforming the previous best method by 14% and even outperforming methods that use external tools by 2.5%. Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource-intensive and requires additional data and expert human annotations. Comparing to in-context learning methods, we also demonstrate that our method is more efficient in terms of the number of steps taken to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
@@ 2 Background
AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018). With the capabilities of LLMs improving, recent work has combined LLMs and RL, training additional policies or value functions to make predictions (Li et al., 2022; Nottingham et al., 2023).
@@ 2.1 In-context learning approaches
Huang et al. (2022a, b) and Singh et al. (2022) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
ReAct Yao et al. (2023b) took this work further by combining ‘acting‘ Huang et al. (2022a) and ‘chain-of-thought‘ Wei et al. (2023). ReAct establishes state-of-the-art for for approaches based only on in-context learning and while it is a very scalable method, the performance (i.e. success rate) is still limited.
ExpeL Zhao et al. (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training. At inference time, they look up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to the fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference. They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few-shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
@@ 2.1.1 In-context learning and additional tools
The current state-of-the-art for in-context learning-based approaches in combination with additional tools is ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner Sun et al. (2023). AdaPlanner introduces a code-based prompt Li et al. (2023a). They use code-execution as an additional tool, by executing LLM generated code and feeding the results into the next LLM generation. The shortcoming of AdaPlanner is that it requires very complex human-crafted prompts that are hard to scale to new environments and it requires the additional step of code-execution.
@@ 2.2 State tracking in LLM-based agents
Concurrent work to ours, AutoGuide Fu et al. (2024), uses ReAct as the base agent as well as additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent; they then use a type of retrieval augmented generation (RAG) process to guide the decision making process. They embed the current observation as a summary (which they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action. Using training data and LLM-based RAG, they achieve 79.1% on top of a ReAct agent on Alfworld. Their training and RAG approach could be used complementary to our StateAct LLM-agent [^2] However, the cost of using RAG at every generation step could be significantly higher than our method..
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task without training data. Their method differs to ours two-fold. Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and, finally, a prediction of a program). Secondly, their system involves execution of actual programs. Our method, on the other hand, requires a straightforward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action. Additionally, we do not require program execution.
@@ 2.2.1 Fine-tuning approaches
2.4 Joined rule and LLM-based agents
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘expert’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al. (2021)).
@@ 2.5 Approaches with multiple trials
Works such as Reflexion Shinn et al. (2023) or work by Sun et al. (2023); Yang et al. (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score. While the score necessarily goes up (as only the failed environments are re-run), we do not dive into this as it does not compare to the actual agent performance; additionally, ‘reflection‘ can be used on top of most LLM-based agents (e.g. Reflexion uses ReAct as the base agent).
@@ 3 Method
StateAct is an LLM-based AI agent that works on top of pre-trained large language models. StateAct takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current interaction trace up to the current step in the environment, see Figure 1(b). An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM. In the case of StateAct, the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The action is then extracted and passed to the environment to produce the next observation, see Figure 2. For example, a sample of the trace for step n in Alfworld is:
@@
In the above example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 1, you see a tomato 1’.
@@ 3.1 Goal-reminders and state-tracking using LLMs
In order to make StateAct more precise, we introduce goal and state tracking. We denote $\pi$ as the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
@@
Where $a_{t}$ is the action produced by the agent at time step t and $o_{t}$ is the observation produced by the environment at time step t after receiving action $a_{t}$ as input. Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
For our case, we need to enhance the policy to incorporate the ‘state’. Similar to previous work Yao et al. (2023b), we introduce the context vector, $c_{t}$. The context vector contains the action, as well as the additional predictions of the agent , i.e. $c_{t}=(g_{0},s_{t},r_{t},a_{t})$. Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before. The new policy $\pi$ then becomes:
@@
In our case, the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
@@ 4 Experimental setup
Our aim is to study the long-range acting, planning and reasoning capabilities of LLM-based agents. To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models. Specifically, we use Alfworld Shridhar et al. (2021), a household robotics environment, and Webshop Yao et al. (2023a), an online shopping environment, as our simulated environments. In line with previous work, we use OpenAI’s [^3] https://openai.com, last accessed June 2024. gpt-3.5-turbo-1106. We, also, evaluate Alfworld on additional LLMs and validate our results.
@@ 4.1 Alfworld
Alfworld Shridhar et al. (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3. Alfworld has a total of 134 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step, the environment provides either a textual observation or feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019), see Appendix A for a complete list of the commands and further details on the environments.
@@ 4.1.1 Alfworld correction
In our research, we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way. Using only “in” or only “on” produces a failed command. We observed this issue with LLMs in this environment and we propose a simple fix for it. We map: 1. “put <object> in <place>” and 2. “put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
Methods such as AdaPlanner Sun et al. (2023) have avoided this issue because they use code-based prompts and regex parsers. However, methods such as ReAct Yao et al. (2023b) and ExpeL Zhao et al. (2023) have been affected by it therefore lowering their potential performance. We report results for ReAct and StateAct both with and without these corrections.
@@ 4.2 Webshop
Figure 4: An textual interaction in Webshop (right) and correspondingwebsite rendering (left).
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results, select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product, see Appendix B for more details. In total, the test set consists of 500 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of 15 steps and two commands: 1. search[<query>], 2. click[<button>].
@@ 4.3 In context learning
Since ReAct Yao et al. (2023b) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024), we use the same few-shot ‘interaction traces’ as ReAct. The main reason is to have a fair comparison and to isolate additional effects - such as performance change - from different in-context examples.
In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens). For our study, we reuse the observations, thoughts and actions, and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well. We release all our annotations with our code release. In comparison, AdaPlanner Sun et al. (2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average.
We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop in all our experiments.